OpenStudy (anonymous):

how to simplify 4 1/3 time 2/5=

OpenStudy (anonymous):

1 11/15(?) I'm not sure I understand this..

OpenStudy (jdoe0001):

\(\bf 4\cfrac{1}{3} \times \cfrac{2}{5}\implies \cfrac{4\times 3+1}{3}\times \cfrac{2}{5}\implies \cfrac{13}{3}\times \cfrac{2}{5}\implies \cfrac{13\times 2}{3\times 5}\)